A rectangular label on a book has an area of 44^cm2 and a perimeter of 27cm.Find the dimensions of the label.If somebody could help me with thes thank you a lot.
Answer by checkley75(3666)

27=2L+2W & L*W=44 L=44/W NOW SUBSTITUTE (44/W) FOR L IN THE FIRST EQUATION & SOLVE FOR W.
27=2(44/W)+2W
27=88/W+2W
27=(88+2W^2)/W
27W=88+2W^2
2W^2-27W+88=0
(2W-11)(W-8)=0
2W-11=0
2W=11
W=11/2
W=5.5 ANSWER.
W-8=0
W=8 ANSWER.
27=2L+2*5.5
27=2L-11
2L=27-11
2L=16
L=8 ANSWER WHEN W=5.5
L=5.5 WHEN W=8.
PROOF
44=8*5.5
44=44